# Learning to accelerate distributed ADMM using graph neural networks

### v1

@@ 4.1 Learning ADMM via message-passing networks
The distributed optimization problem can be directly encoded as the input to a GNN. The underlying graph structure is naturally given by the communication network $\mathcal{G}$ from the problem formulation (2). The node features of the graph consist of the three ADMM iterates $(x_{i},y_{i},\lambda_{i})$. Furthermore, the non-negative edge weights $e_{ij}>0$ of the graph represent the weighted connectivity, that determine the strength of the connection between two nodes in the graph. To obtain a valid communication matrix from the edge features, we construct the weighted Laplacian $P=\tilde{D}-E$ where $\tilde{D}$ is a diagonal $m\times m$ matrix of the weighted degree for each node $\tilde{D}_{ii}=\sum_{j\in\mathcal{N}(i)}e_{ji}$ and $E$ the symmetric matrix representation of the graph connectivity.
@@
We now explain our reformulation of an ADMM iteration in form of two message-passing steps. Recall that in the MPNN framework in (6), the node update function only has access to the aggregation of the incoming messages and not the individual messages, as required in the original ADMM update step (4). To overcome this difficulty and define a valid update function, we rewrite (4) equivalently to
@@
In the rewritten form, $(\bar{\lambda}_{\rightarrow i}^{k},\bar{y}_{\rightarrow i}^{k})$ is the summation (agg) of all to node $i$ incoming messages that we define by $(P_{ji}\lambda_{j}^{k},P_{ji}y_{j}^{k})$. Hence, in accordance with the MPNN framework, the refined update function in equation (7) now only depends on the aggregated features and the nodes’ local information that is encoded in the node features. For details of the derivation, we refer to Appendix D.
The updates for $y$ and $\lambda$ in equation (5) can be carried out jointly within an second message-passing step, and their translation is straightforward. It suffices to define the message between two adjacent nodes as $P_{ij}x_{j}^{k+1}$ and replace the respective part in (5) with the aggregate of these messages.
Together, this yields a two-step message-passing network that we visualize in Figure 1 and summarize in Appendix D in Algorithm 2.2. We highlight the step for the $x$-update by a red box, and the $y$- and $\lambda$-update by a dashed blue box. Due to the equivalence to the original iteration in (4) and (5), the network’s output coincides with the updated iterates of a single distributed ADMM iteration.
@@
The message-passing steps derived in the previous paragraph rigidly reproduce the decentralized distributed ADMM iterations, without any added flexibility through machine learning. In other words, there are no learnable parameters that can be utilized to improve the performance of the iterations. In the following, we present several ways to introduce learnable parameters in the message-passing steps that still maintain convergence under assumptions stated in Section 3.1. These approaches correspond to different levels of tasks for the GNN:
@@
  • Edge-level task: For each edge in the network, we can learn the positive edge weight $e_{ij}$ between two connected agents. Based on the edge weights predicted before the first ADMM iteration, we use the corresponding weighted Laplacian $P$ as a fixed communication matrix across all steps.
For the first two tasks, the step size is predicted using a small neural network that receives as input the optimization variables of the previous iteration, consisting of the primal variable $x$, and the variables $y$ and $\lambda$ for each node. Further, we append the aggregated messages and the total number of nodes in the graph to the input. We apply an instance normalization layer before passing the input through the network to ensure numerical stability [39]. At the beginning of the algorithm, we can predict and change the step size in every iteration, resulting in a separate MLP for every change. However, to ensure convergence, we fix the predicted step size after a predefined number of $L$ iterations. [^2] In practice, we often choose $L=K$.
For the predicted edge weights, we use the local degree profile [10], describing the local connectivity of the graph, as an input to an MLP. The feature representations of the two nodes get stacked and passed through the learned network to predict a positive edge weight that preconditions the ADMM iterations. To ensure invariance to the node order and symmetrize the edge weight graph, we sum the predicted edge weights for each directed edge. To ensure the positivity of the learned algorithm hyperparameters $e_{ij}$ and $\alpha$, we use the softplus activation function on the network output.
@@ 4.2 Network training
In real-world applications, distributed optimization problems need to be solved repeatedly for similar communication networks $\mathcal{G}$ and similar local objectives $f_{i}$, providing potential for problem-class specific acceleration. We assume access to a dataset $\mathcal{D}$ consisting of a sequence of problem instances that share an underlying structure to train our model to improve the average performance across the dataset.
@@
We apply a technique called algorithm unrolling [12], where the algorithm is trained on a fixed number of iterations to maximize performance within this iteration budget. In our case, we construct a GNN, consisting of $K$ distributed ADMM iterations via the message-passing scheme defined in the previous section. Thus, in total we use $2K$ message passing steps for the GNN implementation.
Depending on the chosen learnable components, several MLPs are incorporated into the GNN to influence the update functions. The parameters of all the neural networks are denoted by $\theta$. The final output contains the approximate solution to the optimization problem (2) which now depends on these parameters. The model is then trained using standard gradient-based learning methods and learns how to choose the learnable components (by adjusting $\theta$) so that a suitable loss function is minimized.
Optimization subproblem
Computing the first part can be achieved in closed form by the means of implicit differentiation [2]. However, implicit differentiation can in practice be costly and requires additional overhead. Instead, we solve the subproblem (7) approximately by unrolling an iterative solver [^3] Here, we unroll the iterative solver within each unrolled iteration of ADMM. or use the closed-form solution to (7) if it exists. In both cases, we compute its gradient using standard automatic differentiation techniques. Along the same lines, the partial derivative of the optimization problem (4) with respect to the learned edge weights $e_{ij}$ is computed. The remaining update steps can be differentiated using standard automatic differentiation tools provided by JAX [9].
@@
The loss function is designed such that it evaluates the GNN outcome $x^{K}=[x_{i}^{K}]_{i=1}^{m}$ only after the $K$ unrolled iterations of ADMM. Thus, we do not care if the learned ADMM iterations perform worse than the baseline method after $k<K$ steps. For practical problems, however, the performance for more iterations of ADMM is also evaluated during testing.
To train our model, we utilize the dataset $\mathcal{D}$. Each problem instance from $\mathcal{D}$ consists of a graph $\mathcal{G}$ and the parameterization of the per-node local objective functions $f_{i}$. Additionally, we have access to the true minimizer $x^{\star}_{d}$ that solves problem (2) for instance $d\in\mathcal{D}$ and can be computed offline efficiently as the problem is convex. Further, we precompute $\hat{x}^{K}_{d,i}]_{i=1}^{m}$, the approximate solution obtained by running the baseline distributed ADMM iteration with the default step size and communication matrix. This is used to normalize the loss function in (8) and avoid over-weighting more involved instances in the training data. We train the network to minimize the empirical risk of the training data by minimizing the loss function
@@
using stochastic gradient-based optimization and automatic differentiation. Here, $m_{d}$ is the number of agents in the problem instance $d\in\mathcal{D}$, and $\varepsilon$ is a small positive constant that prevents division by zero.
This loss function is regression-based as it relies on the true solution to the optimization problem (2) that can be computed apriori and is part of $\mathcal{D}$ [1]. Further, the loss function only cares about minimizing the distance to $x^{\star}_{d}$ without directly training the network to fulfill the consensus constraint (2b) or to minimize the global objective (2a). The denominator does not depend on the network parameters but is used to normalize the loss functions across problem instances. Moreover, the normalized loss allows a natural interpretation of the network performance as it measures the network outcome relative to the baseline method with fixed hyperparameters.
@@ 5.1 Problem setting
We present and evaluate our method on two different use-cases of the distributed ADMM algorithm which are inspired by real-world problems. More details about the problem generation for the problems can be found in Appendix B.
@@
The first problem we consider is network consensus where all agents have some local information $b_{i}$ and the goal for the agents is to find the mean of all information across the network [43]. Despite its simplicity, this problem is a classic benchmark that highlights some of the important design choices in distributed optimization. The objective function for each agent $i$ in the network can be encoded as:
$f_{i}(x_{i})=\|x_{i}-b_{i}\|^{2}.$ (9)
@@
For the second use-case we consider an application area inspired by machine learning where we want to solve a distributed least squares problem [41]. Here, each node has access to only a part of the data given by inputs $B_{i}$ and corresponding labels $b_{i}$ and the goal of the agents is to collaboratively find the least-squares solution xx. [^4] In the machine learning literature, the data is usually referred to as $(x,y)$ pairs and the parameters of the linear regression model as $\theta$. In order to be consistent with the ADMM framework presented above, we use $x$ as the regression parameters and $(B_{i},b_{i})$ as the data. Thus, the local objective is:
$f_{i}(x_{i})=\|B_{i}x_{i}-b_{i}\|^{2}.$ (10)
@@ 5.2 Results
We train our model for $K=10$ unrolling steps to minimize the loss function (8). Additional details about the model training can be found in Appendix C. In Table 1, we show the results of different learned methods in terms of their ability to minimize the loss function (8) and achieve consensus among the iterates across the network.
In the middle column, we show the results obtained for $K=10$ iterations which is the value used for the loss function in (8). We can see that all learned methods outperform the baseline, achieving a lower error across all learned methods in both experimental settings. This also holds for consensus among the iterates, even though the loss only influences this indirectly. Overall, the combined method clearly performs best in both metrics.
Table 1: Comparison of different methods of the ADMM algorithm after a fixed number of K steps. Numbers reported are the average across the test data. The error measures the average distance of the local solutions to the global minimizer $\texttt{avg}\left(\|x_{i}^{K}-x^{\star}\|^{2}_{2}\right)$ and the consensus measures the degree to which the constraint (2b) is satisfied as $\texttt{avg}\left(\|x_{i}^{K}-\bar{x}^{K}\|\right)$, where $\bar{x}^{K}$ is the average solution at step $K$. The baseline method does not use learning but uses the default values $\alpha=1,e_{ij}=1$ for all iterations instead. In the combined version, we learn both individual step-size parameters $\alpha_{i}$ and the edge weights simultaneously. Lower is better for all metrics. Best is bold, second base underlined.
k=5 K=10 (train) k=20
Consensus Error Consensus Error Consensus Error Consensus
Baseline 19.07 12.23 9.17 7.10 2.33 2.13
Global $\alpha$ 18.74 17.21 4.19 3.65 3.46 3.37
Local $\alpha_{i}$ 14.50 12.26 3.05 2.82 1.17 1.14
Weight $e_{ij}$ 17.98 10.96 7.39 5.47 1.52 1.40
Combined 13.39 8.05 1.96 1.76 0.76 0.65
Least-squares Error Consensus Error Consensus Error Consensus
Baseline 70.86 10.29 53.23 7.77 30.09 2.44
Global $\alpha$ 61.32 23.41 27.12 9.21 14.93 3.48
Local $\alpha_{i}$ 56.80 14.66 23.79 7.37 12.96 1.84
Weight $e_{ij}$ 67.83 14.19 43.99 8.20 19.45 2.51
Combined 58.97 8.83 18.24 5.42 11.79 0.95
@@
In the leftmost column of Table 1, we compare the results for $k=5$ steps. We can see that generally, the gap between the baseline method and the learned methods is not as big as for the training objective. However, already after $5$ iterations, the learned methods outperform the baseline in most cases, even though the training only evaluates the performance after $K=10$ steps.
In the rightmost column, we show the performance of the (learned) ADMM algorithm after $k=20$ iterations. For the additional iterations, we use a constant step size $\alpha=1$ as in the baseline. This allows us to guarantee the convergence of the ADMM iterations by fixing the step size for all remaining iterations. Moreover, we apply the same learned weighted Laplacian matrix in the additional iterations. Also in this situation, apart from the global step size in the consensus problem, we observe advances compared to the baseline. Unsurprisingly, the combined method shows the strongest improvements.
In summary, although our method is trained to perform well only after the fixed $K=10$ iterations, both earlier and later iterates still outperform the baseline method in our experiments.
@@
In addition to the comparisons in the previous paragraph, we also evaluate the obtained results in terms of the objective function value. We measure this by the relative objective at iteration $k$, which is the difference in the objective values normalized by the minimal value:
$\frac{\left|f(x^{k})-f(x^{\star})\right|}{|f(x^{\star})|}\,.$ (11)
In Figure 2, the relative objective is plotted for $k=20$ iterations for the different methods. We can clearly see that the performance of the learned methods is best after the $K=10$ unrolling steps for training. The model using a global $\alpha$ even shows a large initial increase in the objective when additional iterations are executed. Further, we can see that most methods initially perform worse than the baseline method but performance improves with more iterations.
One exception is the learned communication matrix for the least-squares problem which leads to a higher objective value for most of the iterations even though the distance to the minimizer gets reduced as we have seen in Table 1. However, when combining the communication matrix with the learned step size, the performance significantly improves also in terms the objective value showing that the preconditioning effect still benefits the ADMM convergence.

### v2

@@ 4.1 Learning ADMM via message-passing networks
The distributed optimization problem can be directly encoded as the input to a GNN. The underlying graph structure is naturally given by the communication network $\mathcal{G}$ from the problem formulation (1). The node features of the graph consist of the three ADMM iterates $(x_{i},y_{i},\lambda_{i})$. Furthermore, the non-negative edge weights $e_{ij}>0$ of the graph represent the weighted connectivity, that determine the strength of the connection between two nodes in the graph. To obtain a valid communication matrix from the edge features, we construct the weighted Laplacian $P=\tilde{D}-E$ where $\tilde{D}$ is a diagonal $m\times m$ matrix of the weighted degree for each node $\tilde{D}_{ii}=\sum_{j\in\mathcal{N}(i)}e_{ji}$ and $E$ is the symmetric matrix representation of the graph connectivity.
@@
We now reformulate the ADMM iteration in form of two message-passing steps. Recall that in the MPNN framework in (5), the update function only has access to the aggregation of the incoming messages and not the individual messages, as required in the original ADMM update step (3). To define a valid update function, we rewrite (3) equivalently to
@@
In the rewritten form, $(\bar{\lambda}_{\rightarrow i}^{k},\bar{y}_{\rightarrow i}^{k})$ is the summation (agg) of all incoming messages to node $i$ that we define by $(P_{ji}\lambda_{j}^{k},P_{ji}y_{j}^{k})$. Hence, in accordance with the MPNN framework, the refined update function in equation  (6) now only depends on the aggregated features and the node’s local information that is encoded in the node features. For details of the derivation, we refer to Appendix E. The updates for $y$ and $\lambda$ in equation (4) can be carried out jointly within a second message-passing step, and their translation is straightforward. It suffices to define the message between two adjacent nodes as $P_{ij}x_{j}^{k+1}$ and replace the respective part in (4) with the aggregate of these messages.
Together, this yields a two-step message-passing network that we visualize in Figure 1. Due to the equivalence to the original iteration in (3) and (4), the network’s output coincides with the updated iterates of a single distributed ADMM iteration.
@@
The message-passing steps derived in the previous paragraph rigidly reproduce the decentralized distributed ADMM iterations, without any learnable parameters that could improve performance. In the following, we present several ways to introduce learnable parameters in the message-passing steps that still maintain convergence under assumptions stated in Section 3.1. These approaches correspond to different levels of tasks for the GNN.
@@
•  Edge-level task: For each edge in the network, we learn the positive edge weight $e_{ij}$ between two connected agents. Based on the edge weights predicted before the first ADMM iteration, we use the corresponding weighted Laplacian $P$ as a fixed communication matrix for all steps.
For the first two tasks, the step size is predicted for each node individually using a neural network that receives as input the corresponding node features $(x_{i},y_{i},\lambda_{i})$. Additionally, we append the aggregated messages and the total number of nodes in the graph to the input. No further information about the local objective $f_{i}$ is used as an input. We apply instance normalization before passing the input through the network (ulyanov2016instance). At the beginning of the algorithm, we can predict and change the step size in every iteration, resulting in a separate MLP for every change. However, to ensure asymptotic convergence, we fix the step size after a predefined number of $L$ iterations. [^2] In practice, we often choose $L=K$.
For the predicted edge weights, we use the local degree profile (cai2018simple), describing the local connectivity of the graph, as an input to an MLP. The feature representations of the two nodes get stacked and passed through the network to predict a positive edge weight that preconditions the ADMM iterations. To ensure invariance to the node order and symmetrize the edge weight graph, we sum the predicted edge weights for each directed edge. To ensure the positivity of the learned algorithm hyperparameters $e_{ij}$ and $\alpha$, we use the softplus activation function.
@@ 4.2 Network training
In real-world applications, distributed optimization problems are solved repeatedly for similar communication networks $\mathcal{G}$ and local objectives $f_{i}$, providing potential for problem-class specific acceleration. We assume access to a dataset $\mathcal{D}$ consisting of problem instances that share an underlying structure to train our model to improve the average performance across the dataset.
@@
We apply a technique called algorithm unrolling (chen2022learning), where the network is trained on a fixed number of iterations to maximize performance within this iteration budget. In our case, we construct a GNN, consisting of $K$ distributed ADMM iterations via the message-passing scheme defined in the previous section. This requires $2K$ message-passing steps for the GNN implementation. Depending on the chosen learnable components, several MLPs are incorporated into the GNN to influence the update functions. The parameters of all the neural networks are denoted by $\theta$. The final output contains the approximate solution to the optimization problem (1) which now depends on these parameters. The model is then trained using standard gradient-based learning methods that adjust the parameters $\theta$ to minimize a suitable loss function.
@@
The loss function is designed such that it evaluates the GNN output $x^{K}=[x_{i}^{K}]_{i=1}^{m}$ only after the final $K$ unrolled iterations of ADMM. Thus, we do not consider the learned ADMM performance after $k<K$ steps, nor how the iterates behave in subsequent iterations after $K$.
To train our model, we utilize a dataset $\mathcal{D}$, where each problem instance consists of a graph $\mathcal{G}$ and the per-node local objective functions $f_{i}$. Additionally, we have access to the true minimizer $x^{\star}_{d}$ that solves problem (1) for instance $d\in\mathcal{D}$ that can be computed offline efficiently since the problem is convex. Further, we precompute $\hat{x}^{K}_{d,i}]_{i=1}^{m}$, the approximate solution obtained by running the distributed ADMM iteration with default hyperparameters. This is used to normalize the loss function in (7) and avoid over-weighting more involved instances in the training data. We train the network to minimize the empirical risk on the training data by minimizing the loss function
@@
using stochastic gradient-based optimization and automatic differentiation. Here, $m_{d}$ is the number of agents in instance $d\in\mathcal{D}$, and $\varepsilon$ is a small positive constant that prevents division by zero.
This loss function is regression-based as it relies on the true solution to the optimization problem (1) that can be computed apriori and is part of $\mathcal{D}$ (amos2023tutorial). As such, the loss penalizes deviations from $x^{\star}_{d}$ without explicitly enforcing the consensus constraint or minimizing the global objective (1). However, consensus is implicitly learned since all agents are trained to reach the same global solution. The denominator does not depend on the network parameters but is used to normalize the loss function across problem instances. Moreover, the normalized loss allows a natural interpretation of the network performance as it measures the network outcome relative to the baseline method with fixed hyperparameters. Using an objective-based loss function instead of (7) requires adding an explicit term for consensus among the agents and a careful design of the normalization.
@@ 5.1 Problem setting
We evaluate our method on two different use cases of distributed ADMM that are inspired by real- world problems. More details about the problem generation can be found in Appendix B.
@@
The first problem we consider is network consensus where all agents have local information $b_{i}\in\mathbb{R}^{n}$ and the goal is to find the mean of all information across the network (zhang2014lasso). Despite its simplicity, this problem is a classic benchmark that highlights important design choices in distributed optimization. The objective function for each agent $i$ in the network is $f_{i}(x_{i})=\|x_{i}-b_{i}\|^{2}$.
@@
For the second use-case, we consider the distributed least-squares problem frequently arising in machine learning (yang2020distributed). Each node has access to only $c$ samples of the data given by $B_{i}\in\mathbb{R}^{c\times n}$ and corresponding labels $b_{i}\in\mathbb{R}^{c}$. The goal is to collaboratively find the least-squares solution with local objective: [^3] In machine learning, the data is usually referred to as $(x,y)$ pairs and the parameters of the model as $\theta$. To be consistent with the ADMM framework above, we use $x\in\mathbb{R}^{n}$ as the regression parameters and $(B_{i},b_{i})$ as the data. $f_{i}(x_{i})=\|B_{i}x_{i}-b_{i}\|^{2}$.
@@ 5.2 Results
We train all our models for $K=10$ unrolling steps to minimize the loss function (7) and compare the results with two baseline methods: (i) a version that uses the best performing $\alpha$ as a fixed global step size, and (ii) an adaptive heuristic that adjusts $\alpha_{i}$ dynamically. Additional details about the model training and baseline selection can be found in Appendix C. In Table 1, we show the results of different learned methods corresponding to individual learning components and a combined version in terms of their ability to minimize the loss function (7) and achieve consensus among the iterates.
In the middle-left column, we show the key results obtained for $K=10$ iterations, which is the assumed computational budget used for the loss function in (7). We can see that all learned methods outperform the baselines, achieving a lower error across all learned methods in both experimental settings. This also holds for most methods in terms of consensus among the iterates, even though the loss only influences this indirectly. Overall, the combined method where we learn both edge weights and step sizes clearly performs best in both metrics.
@@
In the leftmost column of Table 1, we compare the results for $5$ steps. Generally, the gap between the baseline methods and the learned methods is not as big as for the training objective. However, already after $5$ iterations, the learned methods outperform the baselines in most cases, even though the training only evaluates the performance after $K=10$ steps.
In the middle-right column, we show the performance of the (learned) ADMM algorithm after $20$ iterations. For the additional iterations, we fix the step size at $\alpha=1$ to guarantee convergence of ADMM and apply the same learned weighted Laplacian matrix. Except for the global step size in the consensus problem, we observe improvements compared to the baselines. Overall, the combined method demonstrates the strongest improvements. Results for $100$ iterations, which is far beyond the training regime, are shown in the rightmost column. Apart from the combined method in the least-squares case, which appears to focus more on reaching consensus, all learned methods improve the baseline metrics. In summary, although our method is trained to perform well only after the trained $10$ iterations, both earlier and later iterates still outperform the baseline methods in our experiments.
@@
In addition, we evaluate the results in terms of the objective function value, measured by the relative objective at iteration $k$, i.e., the difference in objective values $\left|f(x^{k})-f(x^{\star})\right|$ normalized by the minimal value $|f(x^{\star})|$. In Figure 2, the relative objective is plotted for $20$ iterations. We can clearly see that the performance benefit of the learned methods is largest after the $10$ unrolling steps used in training. When additional iterations are executed, the model using a global $\alpha$ shows a large initial increase in the objective, likely due to overfitting on the number of unrolling iterations and freezing the step size. In the least-squares case, we also observe that the method using the learned communication matrix leads to a higher objective value for most of the iterations, even though the distance to the minimizer decreases. However, when combining the communication matrix with the learned step size, the performance significantly improves, beyond simply adding the two individual contributions of these hyperparameters. This shows that preconditioning still benefits the ADMM convergence, especially in synergy with learned step sizes. More results comparing wall-clock times and out-of-distribution generalization of the methods can be found in Appendix D.
Table 1: Comparison of different methods of the ADMM algorithm after a fixed number of $k$ steps. Numbers reported are the average across the $100$ test instances. The error measures the average distance of the local solutions to the global minimizer $\texttt{avg}\left(\|x_{i}^{k}-x^{\star}\|^{2}_{2}\right)$ and the consensus gap measures the degree to which the consensus constraint is violated as $\texttt{avg}\left(\|x_{i}^{k}-\bar{x}^{k}\|\right)$, where $\bar{x}^{k}$ is the average solution at step $k$. We compare with a fixed baseline using the best constant step size and an adaptive method which applies a dynamic selection of node-level step sizes. In the combined version, we learn both individual step size parameters $\alpha_{i}$ and the edge weights simultaneously. Lower is better for all metrics. Best is bold, second best underlined.
$k=5$ $k=10$ (train) $k=20$ $k=100$
Consensus Error Consensus Error Consensus Error Consensus Error Consensus
Fixed 18.54 11.02 8.99 6.48 2.26 2.02 0.144 0.144
Adaptive 19.02 12.45 8.86 6.98 2.13 1.94 0.140 0.139
Global $\alpha$ 18.74 17.21 4.19 3.65 3.46 3.37 0.104 0.104
Local $\alpha_{i}$ 14.50 12.26 3.05 2.82 1.17 1.14 0.112 0.112
Weight $e_{ij}$ 17.98 10.96 7.39 5.47 1.52 1.40 0.036 0.036
Combined 13.39 8.05 1.96 1.76 0.76 0.65 0.001 0.001
Least-squares Error Consensus Error Consensus Error Consensus Error Consensus
Fixed 70.89 10.17 53.35 7.68 30.19 2.43 1.28 0.04
Adaptive 70.81 10.64 52.79 7.95 29.58 2.31 1.25 0.04
Global $\alpha$ 61.32 23.41 27.12 9.21 14.93 3.48 0.67 0.04
Local $\alpha_{i}$ 56.80 14.66 23.79 7.37 12.96 1.84 0.62 0.03
Weight $e_{ij}$ 67.83 14.19 43.99 8.20 19.45 2.51 0.37 0.03
Combined 58.97 8.83 18.24 5.42 11.79 0.95 1.94 0.01